# DeRO: Dead Reckoning Based on Radar Odometry With Accelerometers Aided for Robot Localization

### v1

@@ Abstract
In this paper, we propose a radar odometry structure that directly utilizes radar velocity measurements for dead reckoning while maintaining its ability to update estimations within the Kalman filter framework. Specifically, we employ the Doppler velocity obtained by a 4D Frequency Modulated Continuous Wave (FMCW) radar in conjunction with gyroscope data to calculate poses. This approach helps mitigate high drift resulting from accelerometer biases and double integration. Instead, tilt angles measured by gravitational force are utilized alongside relative distance measurements from radar scan matching for the filter’s measurement update. Additionally, to further enhance the system’s accuracy, we estimate and compensate for the radar velocity scale factor. The performance of the proposed method is verified through five real-world open-source datasets. The results demonstrate that our approach reduces position error by 47% and rotation error by 52% on average compared to the state-of-the-art radar-inertial fusion method in terms of absolute trajectory error.
OPEN SOURCE CODE
The implementation of the method described in this article is available at https://github.com/hoangvietdo/dero.
@@ I INTRODUCTION
To the best of the author’s knowledge, existing literature predominantly treats radar as an auxiliary sensor within a loosely coupled framework. This methodology is commonly referred to as radar inertial odometry (RIO). In this framework, radar measurements typically serve the purpose of updating the state estimation, often implemented through an extended Kalman filter (EKF) [11]. Meanwhile, an Inertial Navigation System (INS) algorithm integrates acceleration to derive velocity, and then integrates again to determine the platform’s position [12]. However, this conventional methodology is notably sensitive to the accuracy of accelerometer bias estimation. Even a modest error in bias estimation, as small as 10%, can lead to large position error [13].
To address these challenges, the dead reckoning (DR) fusion technique has been proposed and implemented [13]. When a velocity sensor such as an odometer is available, it replaces the accelerometer for position estimation. This substitution offers several advantages. Firstly, it completely prevents significant errors caused by accelerometers biases (typically from a low-grade IMU), while also eliminating the need for a second integration step to calculate position estimation. Additionally, when relying on accelerometer measurements, compensating for gravitational forces becomes necessary [12]. Inaccuracies in determining this constant can lead to notable errors due to the effect of double integration. Moreover, conventional INS may include additional attitude errors caused by vibrations, such as those encountered in drones or legged robots. Vibration induces acceleration biases in IMU measurements, whereas it is typically considered as noise in velocity sensors. Consequently, handling noise is generally considered more straightforward than addressing biases.
@@
    A framework of dead reckoning based on radar odometry (DeRO) with accelerometers aided is presented. In this pipeline, the radar’s ego velocity and angular velocity measurement from gyroscope are used for the IEKF’s time update. Simultaneously, the radar’s range measurement and title angles computed by accelerometers contribute to IEKF’s measurement update under stochastic cloning concept.
@@
    We implement the proposed method using C++ in robot operating system 2 (ROS 2) [14] and make our source code openly available for the benefit of the research community.
The rest of this article is organized as follows. In Section II, we review related work in the context of RIO and DR. Definitions on coordinate frames and mathematical notations along with DR-based radar odometry mechanization are given in Section III. We develop the stochastic cloning based indirect EKF (IEKF) starting from the state-space definition in Section IV. Next, Section V describes the real-world experiment results to demonstrate our proposed framework. Finally, Section VI concludes this article.
@@ II-A Dead Reckoning
The DR method has been extensively researched and popularly adopted as a straightforward and computationally effecient solution for odometer/INS/GNSS integrated system in GNSS-denied environments [15, 16, 17]. Simulation and field test results have consistently shown the effectiveness of this approach over the conventional INS-based system [13, 18]. To further improve its robustness, adaptive rules for the Extended Kalman Filter (EKF) were introduced by [19] to handle measurement outliers and dynamic model disturbances. In addition, Takeyama et al. [20] extended the traditional DR-based structure by incorporating a monocular camera to improve heading estimation accuracy.
@@
One thing we notice from these previous researches is that, most of the DR-based application has primarily been applied to land vehicles, with limited extensions to other platforms such as drones. This restriction stems from the availability of odometers in wheeled vehicles, their cost-effectiveness, and they require minimal effort to compute the vehicle’s body velocity. Moreover, it is not straightforward to directly measurement body velocity in other non-wheeled applications. In contrast to previous approaches, our work employs a 4D FMCW radar instead of an odometer, thus overcoming this limitation.
@@ III-B System Overview
The overview of the proposed DeRO is depicted in Fig. 2. In this study, we apply a stochastic cloning based error-state (indirect) EKF (SC-IEKF) to fuse data from an IMU and a 4D FMCW radar. The raw point clouds obtained from radar are first fed into an outlier rejection algorithm. Here, we employ a simple RANSAC-based least square (RANSAC/LSQ) method introduced in [11]. The outputs of this algorithm include the estimated ego velocity $\hat{v}^{r}$ along with its corresponding error covariance $\hat{Q}$, as well as a set of inlier point clouds. These velocity estimates and covariance matrices are used for radar mechanization in conjunction with gyroscope data to compute poses and the filter’s time update. Subsequently, the inlier point cloud set is merged with the inlier set from the $M$ preceding steps to perform scan matching to calculate the relative distance estimate $\hat{p}^{r_{k}}_{r_{k-M}}$ with error covariance $\hat{R}$ between these two radar scans. This information will update the estimation.
It is important to note that we operate under the assumption of a static environment. In other words, throughout the entire process, no moving objects are present within the field of view (FOV) of the radar. Chen et al. [31] have demonstrated that the accuracy of ego velocity estimation from radar degrades rapidly in the presence of moving objects within the scene. However, addressing this issue falls outside the scope of this article.
@@ III-C Radar mechanization
$v^{r}$ $=C_{b}^{r}\bigg{(}C_{n}^{b}v_{b}^{n}+\left[\omega^{b}\right]_{\!\times}p_{r}^{b}\bigg{)}$ (1)
@@ IV-A State Augmentation
with the position $p_{b}^{n}$ and the unit quaternion $q_{b}^{n}\in\mathbb{R}^{4}$ with the norm $\|q_{b}^{n}\|=1$ are both referenced to the global frame. The vector $b_{g}$ represents the gyroscope biases expressed in the body frame, while $s_{r}\in\mathbb{R}^{3}$ denotes the radar scale factor.
@@
Therefore, the error-state vector of INS is given by
@@ IV-B System Model
$\delta\dot{p}_{b}^{n}$ $=\delta v_{b}^{n}$
$=-\hat{C}_{b}^{n}C_{r}^{b}\textup{diag}(\bar{v}^{r})\textup{diag}(\hat{s}_{r})^{-1}\delta s_{r}+\hat{C}_{b}^{n}C_{r}^{b}n_{r}$
$\mathrel{\phantom{=}}-\left[\hat{C}_{b}^{n}\left[\hat{\omega}^{b}\right]_{\!\times}p_{r}^{b}-\hat{C}_{b}^{n}C_{r}^{b}\hat{v}^{r}\right]_{\!\times}\delta\Psi_{b}^{n}-\hat{C}_{b}^{n}\left[p_{r}^{b}\right]_{\!\times}\delta b_{g}.$
$\mathrel{\phantom{=}}+\hat{C}_{b}^{n}\left[p_{r}^{b}\right]_{\!\times}n_{g}$ (17)
Here, we use the fact that the true value of the scale factor is $s=1$. Additionally, we remove the hat symbol above the IMU-Radar calibration parameters $C_{r}^{b}$ and $p_{r}^{b}$ based on the assumption established in Section III-C. Thus, they become constant.
@@
In this context, the process noise vector $w_{k-1}$ and its corresponding matrix $G_{k-1}$ (for brevity, only non zero elements of $G_{k-1}$ are provided) are respectively defined as
$w_{k-1}$ $=\begin{bmatrix}n_{r}^{\top}&n_{g}^{\top}&n_{b_{g}}^{\top}&n_{s_{r}}^{\top}\end{bmatrix}^{\top}$ (19b)
$G_{k-1}(0:2,0:2)$ $=\hat{C}_{b}^{n}C_{r}^{b},\,\,G_{k-1}(0:2,3:5)=\hat{C}_{b}^{n}\left[p_{r}^{b}\right]_{\!\times}$ (19c)
$G_{k-1}(3:5,3:5)$ $=-\hat{C}_{b}^{n},\,\,G_{k-1}(6:11,6:11)=I_{6}$ (19d)
The transition matrix $\Phi_{k-1}$ is discretized from
@@
where $T$ represents the sampling period, and the dynamic matrix $F\in\mathbb{R}^{12\times 12}$ is defined (for brevity, only non zero elements of $F$ are provided)
@@
$F(0:2,9:11)$ $=-\hat{C}_{b}^{n}C_{r}^{b}\textup{diag}(\bar{v}^{r})\textup{diag}(\hat{s})^{-1}$ (21c)
@@
where $\Xi=\Upsilon_{k-1}P_{k-1}^{+}\Upsilon_{k-1}^{\top}+G_{k-1}Q_{d,k-1}G_{k-1}^{\top}$ with the discretized process noise covariance $Q_{d,k-1}=T\textrm{diag}\big{(}\sigma_{g}^{2},\sigma_{r}^{2},\sigma_{b_{g}}^{2},\sigma_{s_{r}}^{2}\big{)}$ and $P_{k-1}^{+}$ is the a posteriori error covariance of the DR error-state $\delta x_{dr,k-1}$. Note that $\Upsilon=\Pi_{i=1}^{M}\Phi_{k-i}$ accumulates over time while awaiting the measurement update step.
@@ IV-C1 Radar range measurements
$\breve{H}_{r,k}=C_{b}^{r}\hat{C}_{n}^{b_{k-M}}\begin{bmatrix}I_{3}&0_{3\times 9}&-I_{3}&\left[\hat{p}_{b_{k-M}}^{n}-\hat{p}_{b_{k}}^{n}\right]_{\!\times}\end{bmatrix},$ (31)
and $n_{p_{r}}\thicksim{\mathcal{N}}\big{(}0_{3\times 1},\sigma^{2}_{p_{r}}I_{3}\big{)}$. The process of deriving the aforementioned matrix can be referenced from [28], with extensions provided for 3D cases.
Unlike in [29], where each target’s distance measurement is directly used, we employ the entire scan for matching to determine the distance between frames. This strategy leads to a significant reduction in computational cost.
@@ IV-C2 Tilt angles
It is well known that roll and pitch angles of the body frame can be computed from accelerometer readings $\bar{f}^{b}$ as [1]
$\bar{\phi}_{a}=\textrm{atan2}\big{(}\dfrac{-\bar{f}_{y}}{-\bar{f}_{z}}\big{)},\quad\bar{\theta}_{a}=\textrm{atan2}\big{(}\dfrac{\bar{f}_{x}}{\sqrt{\bar{f}_{y}^{2}+\bar{f}_{z}^{2}}}\big{)},$ (32)
@@
$z_{a,k}$ $=\begin{bmatrix}\hat{\phi}\\ \hat{\theta}\end{bmatrix}-\begin{bmatrix}\bar{\phi}_{a}\\ \bar{\theta}_{a}\end{bmatrix}=H_{a,k}\delta\Psi_{b,k}^{n}+n_{a}$ (37)
@@
Here, the estimate Euler angle $\hat{\Psi}_{b}^{n}=\begin{bmatrix}\hat{\phi}&\hat{\theta}&\hat{\psi}\end{bmatrix}^{\top}$ can be found by transforming the estimate quaternion $\hat{q}_{b}^{n}$ and $n_{a}\thicksim{\mathcal{N}}\big{(}0_{3\times 1},\sigma^{2}_{a}I_{3}\big{)}$. Furthermore, given that velocity information is available from radar measurements, we compensate for the linear acceleration derived from accelerometer output. This compensation enhances the quality of tilt angle information. Particularly,
$\hat{f}^{b}=\bar{f}^{b}-C_{n}^{b}\dfrac{\big{(}v_{b_{k}}^{n}-v_{b_{k-1}}^{n}\big{)}}{T}.$ (41)
Note that we assume accelerometer biases have been calibrated beforehand by another algorithm (e.g., coarse alignment). To further prevent adverse effects of linear acceleration and biases, we adopt the adaptive rule proposed by Park et al. [13]. More specifically, we compare the difference $\|\hat{f}^{b}\|-\|g\|$, where $g$ represents the gravitational force vector, with a predetermined constant threshold $\gamma$. If this difference exceeds $\gamma$, we increment the value of $\sigma^{2}_{a}$. This approach has been effectively employed in [21].

### v2

@@ Abstract
In this paper, we propose a radar odometry structure that directly utilizes radar velocity measurements for dead reckoning while maintaining its ability to update estimations within the Kalman filter framework. Specifically, we employ the Doppler velocity obtained by a 4D Frequency Modulated Continuous Wave (FMCW) radar in conjunction with gyroscope data to calculate poses. This approach helps mitigate high drift resulting from accelerometer biases and double integration. Instead, tilt angles measured by gravitational force are utilized alongside relative distance measurements from radar scan matching for the filter’s measurement update. Additionally, to further enhance the system’s accuracy, we estimate and compensate for the radar velocity scale factor. The performance of the proposed method is verified through five real-world open-source datasets. The results demonstrate that our approach reduces position error by 62% and rotation error by 66% on average compared to the state-of-the-art radar-inertial fusion method in terms of absolute trajectory error.
@@ I INTRODUCTION
To the best of the author’s knowledge, existing literature predominantly treats radar as an auxiliary sensor within a loosely coupled framework. This methodology is commonly referred to as radar inertial odometry (RIO). In this framework, radar measurements typically serve the purpose of updating the state estimation, often implemented through an extended Kalman filter (EKF) [11]. Meanwhile, an inertial navigation system (INS) algorithm integrates acceleration to derive velocity, and then integrates again to determine the platform’s position [12, Ch. 11]. However, this conventional methodology is notably sensitive to the accuracy of accelerometer bias estimation. Even a modest error in bias estimation, as small as 10%, can result in significant position errors [13].
To address these challenges, the dead reckoning (DR) fusion technique has been proposed and implemented [13]. When a velocity sensor such as an odometer is available, it replaces the accelerometer for position estimation. This substitution offers several advantages. Firstly, it completely prevents significant errors caused by accelerometers biases (typically from a low-grade IMU), while also eliminating the need for a second integration step to calculate position estimation. Additionally, when relying on accelerometer measurements, compensating for gravitational forces becomes necessary [12]. Inaccuracies in determining this constant can lead to notable errors due to the effect of double integration. Moreover, conventional INS may include additional attitude errors caused by vibrations, such as those encountered in drones or legged robots. Vibration induces acceleration biases in IMU measurements, whereas it is typically considered as noise in velocity sensors. Generally, handling noise is considered more straightforward than addressing biases.
@@
    A framework of Dead reckoning based on Radar Odometry (DeRO) with accelerometers aided is presented. In this pipeline, the radar’s ego velocity and angular velocity measurement from gyroscope are used for the filter’s time update. Simultaneously, the radar’s range measurement and tilt angles computed by accelerometers contribute to the estimation correction under stochastic cloning concept.
@@
    We implement the proposed method using C++ in robot operating system 2 (ROS 2) [14] and make our source code openly available [^1] https://github.com/hoangvietdo/dero. for the benefit of the research community.
The rest of this article is organized as follows. In Section II, we review related work in the context of RIO and DR. Definitions on coordinate frames and mathematical notations along with DR-based radar odometry mechanization are given in Section III. We develop the stochastic cloning based indirect EKF starting from the state-space definition in Section IV. Next, Section V describes the real-world experiment results to demonstrate our proposed framework. Finally, Section VI concludes this article.
@@ II-A Dead Reckoning
The dead reckoning method has been extensively researched and popularly adopted as a straightforward and computationally effecient solution for odometer/INS/GNSS integrated system in GNSS-denied environments [15, 16, 17]. Simulation and field test results have consistently shown the effectiveness of this approach over the conventional INS-based system [13, 18]. To further improve its robustness, adaptive rules for the EKF were introduced by [19] to handle measurement outliers and dynamic model disturbances. In addition, Takeyama et al. [20] extended the traditional DR-based structure by incorporating a monocular camera to improve heading estimation accuracy.
@@
One thing we notice from these previous researches is that, most of the DR-based application has primarily been applied to land vehicles, with limited extensions to other platforms such as drones. This restriction stems from the availability of odometers in wheeled vehicles, their cost-effectiveness, and they require minimal effort to compute the vehicle’s body velocity. Moreover, it is not straightforward to directly measure body velocity in other non-wheeled applications. In contrast to previous approaches, our work employs a 4D FMCW radar instead of an odometer, thus overcoming this limitation.
@@ III-B System Overview
The overview of the proposed DeRO is depicted in Fig. 2. In this study, we apply a stochastic cloning based error-state (indirect) EKF (SC-IEKF) to fuse data from an IMU and a 4D FMCW radar. The raw point clouds obtained from radar are first fed into an outlier rejection algorithm. Here, we employ a simple RANSAC-based least square (RANSAC/LSQ) method introduced in [11]. The outputs of this algorithm include the estimated ego velocity $\hat{v}^{r}$ along with its corresponding error covariance $\hat{Q}$, as well as a set of inlier point clouds. These velocity estimates and covariance matrices are used for radar mechanization in conjunction with gyroscope data to compute poses and the filter’s time update. Subsequently, the inlier point cloud set is merged with the inlier set from the one at $M$ preceding steps to perform scan matching to calculate the relative distance estimate $\hat{p}^{r_{k}}_{r_{k-M}}$ with error covariance $\hat{R}$ between these two radar scans. This information will update the estimation.
It is important to note that we operate under the assumption of a static environment. Chen et al. [31] have demonstrated that the accuracy of ego velocity estimation from radar degrades rapidly in the presence of moving objects within the scene. However, in the case that the number of static object dominates the moving object, then RANSAC might be sufficient to remove the influence of moving objects. Thus, our assumption can be relative relaxed. Addressing scenarios with a high number of moving objects falls outside the scope of this article.
@@ III-C Radar mechanization
$v^{r}$ $=C_{b}^{r}\bigg{(}C_{n}^{b}v_{b}^{n}+\left[\omega^{b}\right]_{\!\times}p_{r}^{b}\bigg{)},$ (1)
@@ IV STOCHASTIC CLONING INDIRECT EXTENDED KALMAN FILTER
The fundamental concept of an SC-based filter is to consider the correlation between the current state estimate and the previous state estimate. This consideration is necessary when previous estimates are used to calculate essential information for filtering purposes, such as estimating the measurement residual in an EKF. This correlation can be effectively accounted for by augmenting the previous and current estimates into a new single state, allowing the new state error covariance to encapsulate their correlation.
@@ IV-A State Augmentation
where the unit quaternion (4-tuple) $q_{b}^{n}\in\mathbb{R}^{4}$ with the norm constraint $\|q_{b}^{n}\|=1$ is referenced to the global frame. The vector $b_{g}$ represents the gyroscope biases expressed in the body frame, while $s_{r}\in\mathbb{R}^{3}$ denotes the radar scale factor.
@@
Therefore, the DR’s error-state vector is given by
@@ IV-B System Model
$\delta\dot{p}_{b}^{n}$ $=\hat{C}_{b}^{n}C_{r}^{b}\textup{diag}(\bar{v}^{r})\delta s_{r}+\hat{C}_{b}^{n}C_{r}^{b}\textup{diag}(\hat{s}_{r})n_{r}+\hat{C}_{b}^{n}\left[p_{r}^{b}\right]_{\!\times}n_{g}$
$\mathrel{\phantom{=}}-\left[\hat{C}_{b}^{n}\left[\hat{\omega}^{b}\right]_{\!\times}p_{r}^{b}-\hat{C}_{b}^{n}C_{r}^{b}\hat{v}^{r}\right]_{\!\times}\delta\Psi_{b}^{n}-\hat{C}_{b}^{n}\left[p_{r}^{b}\right]_{\!\times}\delta b_{g}.$ (14)
Here, we remove the hat symbol above the IMU-Radar calibration parameters $C_{r}^{b}$ and $p_{r}^{b}$ based on the assumption established in Section III-C that makes them constant.
@@
In this context, the process noise vector $w_{k-1}$ and its corresponding matrix $G_{k-1}$ (for brevity, only non zero elements of $G_{k-1}$ are provided [^2] Using C++ 0-based indexing notations.) are respectively defined as
$w_{k-1}$ $=\begin{bmatrix}n_{r}^{\top}&n_{g}^{\top}&n_{b_{g}}^{\top}&n_{s_{r}}^{\top}\end{bmatrix}^{\top}$ (16a)
$G_{k-1}(0:2,0:2)$ $=\hat{C}_{b}^{n}C_{r}^{b}\textup{diag}(\hat{s}_{r})$ (16b)
$G_{k-1}(0:2,3:5)$ $=\hat{C}_{b}^{n}\left[p_{r}^{b}\right]_{\!\times}$ (16c)
$G_{k-1}(3:5,3:5)$ $=-\hat{C}_{b}^{n}$ (16d)
$G_{k-1}(6:11,6:11)$ $=I_{6}$ (16e)
The transition matrix $\Phi_{k-1}$ is discretized as
@@
where $T$ represents the radar sampling period, and the dynamic matrix $F\in\mathbb{R}^{12\times 12}$ is defined (for brevity, only non-zero elements of $F$ are provided2)
@@
$F(0:2,9:11)$ $=\hat{C}_{b}^{n}C_{r}^{b}\textup{diag}(\bar{v}^{r})$ (18c)
@@
where $\Xi=\Phi_{k-1}P_{k-1}^{+}\Phi_{k-1}^{\top}+G_{k-1}Q_{d,k-1}G_{k-1}^{\top}$ with the discretized process noise covariance $Q_{d,k-1}=T\textrm{diag}\big{(}\sigma_{g}^{2},\sigma_{r}^{2},\sigma_{b_{g}}^{2},\sigma_{s_{r}}^{2}\big{)}$ and $P_{k-1}^{+}$ is the a posteriori error covariance of the DR error-state $\delta x_{dr,k-1}$. Note that $\Upsilon=\Pi_{i=1}^{M}\Phi_{k-i}$ accumulates over time while awaiting the measurement update step.
@@ IV-C1 Radar range measurements
$\breve{H}_{r,k}$ $=C_{b}^{r}\hat{C}_{n}^{b_{k-M}}\begin{bmatrix}I_{3}&\left[\hat{C}_{b_{k}}^{n}p_{r}^{b}\right]_{\!\times}&0_{3\times 6}&-I_{3}&\Lambda\end{bmatrix},$ (23)
$\Lambda=\left[\hat{p}_{b_{k-M}}^{n}-\hat{p}_{b_{k}}^{n}-\hat{C}_{b_{k}}^{n}p_{r}^{b}\right]_{\!\times}$, and $n_{p_{r}}\thicksim{\mathcal{N}}\big{(}0_{3\times 1},\sigma^{2}_{p_{r}}I_{3}\big{)}$.
The process of deriving the aforementioned matrix can be referenced from [28], with extensions provided for 3D cases. Unlike in [29], where each target’s distance measurement is directly used, we employ the entire scan for matching to determine the distance between frames. This strategy leads to a significant reduction in computational cost.
@@ IV-C2 Tilt angles
It is well known that roll and pitch angles of the body frame can be computed from the accelerometer readings $\bar{f}^{b}$ without acceleration as [1]
$\bar{\phi}_{a}=\arctan\bigg{(}\dfrac{-\bar{f}_{y}}{-\bar{f}_{z}}\bigg{)},\,\,\bar{\theta}_{a}=\arctan\Bigg{(}\dfrac{\bar{f}_{x}}{\sqrt{\bar{f}_{y}^{2}+\bar{f}_{z}^{2}}}\Bigg{)},$ (24)
@@
$z_{a,k}$ $=\begin{bmatrix}\hat{\phi}\\ \hat{\theta}\end{bmatrix}-\begin{bmatrix}\bar{\phi}_{a}\\ \bar{\theta}_{a}\end{bmatrix}=H_{a,k}\delta\Psi_{b,k}^{n}+n_{a}$
@@
Here, the estimate Euler angle $\hat{\Psi}_{b}^{n}=\begin{bmatrix}\hat{\phi}&\hat{\theta}&\hat{\psi}\end{bmatrix}^{\top}$ can be found by transforming the estimate quaternion $\hat{q}_{b}^{n}$ and $n_{a}\thicksim{\mathcal{N}}\big{(}0_{2\times 1},\sigma^{2}_{a}I_{2}\big{)}$. Furthermore, given that velocity information is available from radar measurements, we compensate for the linear acceleration and biases derived from accelerometer output. This compensation enhances the quality of tilt angle information. Particularly,
$\hat{f}^{b}=\bar{f}^{b}-\hat{b}_{a}-\hat{C}_{n}^{b}\dfrac{\big{(}\hat{v}_{b_{k}}^{n}-\hat{v}_{b_{k-1}}^{n}\big{)}}{T}.$ (26)
It is important to note that accelerometer sensors are assumed to be pre-calibrated by a separate algorithm (e.g., coarse alignment) to obtain $\hat{b}_{a}$. To further prevent adverse effects of linear acceleration and biases, we adopt the adaptive rule proposed by Park et al. [13]. More specifically, we compare the difference $\|\hat{f}^{b}\|-\|g\|$, where $g$ represents the gravitational force vector, with a predetermined constant threshold $\gamma$. If this difference exceeds $\gamma$, we increment the value of $\sigma^{2}_{a}$. This approach has been effectively employed in [21].